determine the sample size needed to estimate the true proportion of people with a disease to within 5 percentage points with 90% confidence.
Answer #1


Let the population proportion be P=0.5 Q=1-P=1-0.5=0.5 The sample size n = ? The maximum error E = 5% = 0.05 Level of signifi